OpenStudy (k8lyn911):

Calculus III: Find an equation of a sphere if one of its diameters has endpoints (3, 5, 6) and (5, 7, 8).

OpenStudy (misty1212):

HI!! this is not as hard as it looks

OpenStudy (misty1212):

find the center first, the average in each coordinate

OpenStudy (misty1212):

you got that?

OpenStudy (anonymous):

anticipation is killing me...

OpenStudy (k8lyn911):

Sorry! Working through some other homework problems. Which equations do we use for that?

OpenStudy (misty1212):

just like the midpoint in two dimensions take the average in each coordinate

OpenStudy (k8lyn911):

So (x1 - x2, y1 - y2, z1 - z2)?

OpenStudy (misty1212):

so, the average

OpenStudy (misty1212):

add up and divide by two

OpenStudy (misty1212):

"no" i meant just like finding the midpoint of \((3,5)\) and \((7,11)\) in two dimensions add and divide by two in each coordinate in my example the midpoint would be \((5,8)\)

OpenStudy (misty1212):

clear or no?

OpenStudy (k8lyn911):

( (3+5)/2, (5+7)/2, (6+8)/2) = (4, 6, 7)?

OpenStudy (misty1212):

right

OpenStudy (misty1212):

before we continue, do you remember how to find the equation of a circle given two endpoints of the diameter? this is identical, except with one extra coordinate

OpenStudy (misty1212):

the equation is going to be \[(x-4)^2+(y-6)^2+(z-7)^2=r^2\]

OpenStudy (k8lyn911):

Not exactly...

OpenStudy (misty1212):

because the center is \((4,6,7)\)

OpenStudy (misty1212):

equation for a circle with radius \(r\) and center \((h,k)\) is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (misty1212):

this is exactly the same, but with one more coordinate

OpenStudy (k8lyn911):

Oh, so it's like the other problems we've had, but backwards! So we add these values squared and take the square root to find r?

OpenStudy (misty1212):

ok lets go slow

OpenStudy (misty1212):

maybe i am not sure what exactly you mean we got the center, we need the radius

OpenStudy (misty1212):

actually we don't need the radius, we just need the square of the radius, so forget the square root business

OpenStudy (k8lyn911):

Okay.

OpenStudy (misty1212):

same distance formula as with two points, only now each point has three coordinates instead of two

OpenStudy (misty1212):

square of the distance between \((3, 5, 6) \) and \( (4, 6, 7)\) is \[((4-3)^2+(6-5)^2+(7-6)^2\] a pretty easy calculation in this case

OpenStudy (misty1212):

for the actual distance it would be \[\sqrt{(4-3)^2+(6-5)^2+(7-6)^2}\]

OpenStudy (misty1212):

you get 3 pretty much instantly right?

OpenStudy (k8lyn911):

Yes.

OpenStudy (misty1212):

final answer:\[(x-4)^2+(y-6)^2+(z-7)^2=3\]

OpenStudy (k8lyn911):

Oh, I see. Thank you so much for explaining it to me! :)

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]